OpenStudy (anonymous):

find the locations where the tangent line to r=theta is vertical on [0,2pi]

OpenStudy (praxer):

r= what theta, sin theta, cos theta, tan theta. ???????????

OpenStudy (anonymous):

just theta

OpenStudy (anonymous):

the graph looks like a spiral

OpenStudy (freckles):

\[\frac{dy}{dx}=\frac{\frac{dr}{d \theta} \sin(\theta)+r \cos(\theta)}{\frac{dr}{d \theta} \cos(\theta)-rsin(\theta)}\]

OpenStudy (freckles):

so you need to find when the bottom is 0

OpenStudy (anonymous):

thank you

OpenStudy (freckles):

http://tutorial.math.lamar.edu/Classes/CalcII/PolarTangents.aspx good notes to have

OpenStudy (freckles):

honestly I'm having trouble finding where \[\cos(\theta)-\theta \sin(\theta)=0 \]

OpenStudy (freckles):

have you solved it?

OpenStudy (freckles):

http://www.wolframalpha.com/input/?i=cos%28theta%29-theta+*sin%28theta%29%3D0 according to wolfram we should get these but i think we have to use a numerical approach instead of an algebraic one